OpenStudy (anonymous):

Simplify: 7√3-4√6+√48-√54

OpenStudy (dan815):

48=6*8

OpenStudy (dan815):

48=16*3

OpenStudy (dan815):

54=9*6

OpenStudy (dan815):

can u see why that would be helpful

OpenStudy (anonymous):

I am suppose to simplify the entire equation

OpenStudy (anonymous):

Wait i understand what you were trying to say

OpenStudy (dan815):

ok go on

OpenStudy (anonymous):

It would equal 11√6-7√12

OpenStudy (dan815):

not quite

OpenStudy (dan815):